# Calle Nicolás Cabrera 13-15, Cantoblanco, 28049, Madrid, Spain

### v1

@@ 1 Introduction
In the Standard Model (SM), with the mass of the detected Higgs boson measured at about 125 GeV ATLAS:2012yve; CMS:2012qbp, electroweak (EW) symmetry breaking is predicted to proceed via a smooth crossover instead of a strong FOEWPT (SFOEWPT) Kajantie:1996mn. As a consequence, the presence of a SFOEWPT would be a clear indication of physics beyond the Standard Model (BSM). A SFOEWPT can occur even in simple extensions of the SM. For instance, models with additional scalar singlets Profumo:2007wc; Espinosa:2011ax; Cline:2012hg, SU(2) doublets Dorsch:2013wja; Basler:2016obg; Goncalves:2021egx; Biekotter:2021ysx; Biekotter:2022kgf; Biekotter:2023eil and/or triplets Patel:2012pi; Niemi:2018asa, effective theories incorporating higher dimensional operators Grojean:2004xa; Camargo-Molina:2021zgz; Camargo-Molina:2024sde, classically conformal models Nardini:2007me; Konstandin:2011dr, or models with extra dimensions Creminelli:2001th can all exhibit a SFOEWPT. As mentioned above, a sufficiently strong FOEWPT can be probed by future space-based GW observatories. The observation of a GW signal would be very helpful in this context to pinpoint the underlying scalar sector responsible for the FOEWPT.
@@
Irrespective of the specific thermal resummation prescription, the main advantage of the $4D$ approach is that the form of the effective potential at one-loop level is relatively simple to implement, while capturing the leading thermal corrections and providing a qualitatively correct picture of the phase transitions in many extensions of the SM. Recent studies using the $4D$ approach have, for instance, been carried out for the singlet-extended SM Chiang:2017nmu; Chiang:2018gsn; Carena:2019une; Cho:2021itv; Fernandez-Martinez:2022stj; Ellis:2022lft; Zhang:2023mnu; Braathen:2025svl, the Two Higgs doublet model Goncalves:2021egx; Biekotter:2022kgf; Biekotter:2021ysx and the triplet-extended SM Chala:2018opy, taking into account current experimental constraints. In some studies also two-loop contributions have been considered in the $4D$ approach Arnold:1992rz; Espinosa:1996qw; Laine:2000kv; Funakubo:2012qc; Bahl:2024ykv. A key drawback of the $4D$ approach is that it is often affected by the poor behavior of the perturbative expansion at finite temperature, leading to sizable renormalization scale and residual gauge dependencies in predictions for physical observables characterizing the phase transitions. In addition, it suffers from infrared divergences and an unsystematic treatment of thermal fluctuations in the evaluation of transition rates Ekstedt:2020abj; Athron:2022jyi. This leads to large theoretical uncertainties in the predictions for these parameters and quantities that are derived from them, such as GW signals produced during the transitions Croon:2020cgk; Lewicki:2024xan; Gould:2024jjt.
@@
Another problem occurring both in the $4D$ and the $3D$ approaches is the presence of residual gauge dependence in predictions for physical observables related to the phase transition. A proposed solution to this issue is the so-called $\hbar$-expansion Patel:2011th, which systematically expands physical quantities in powers of a loop-counting parameter $\hbar$. This ensures that one obtains gauge-independent results for physical observables order by order in perturbation theory. At the one-loop level in the $4D$ approach, applying the $\hbar$-expansion removes gauge dependence, but leads to results that, in many models, are not sufficiently precise, deviating substantially from more accurate predictions Patel:2011th. Additionally, the $\hbar$-expansion must be implemented consistently with the thermal resummation procedure to obtain a gauge-independent formulation at finite temperature Ekstedt:2018ftj; Athron:2022jyi.
@@
Specifically, we investigate the EW phase transition in the SM extended by a complex gauge singlet scalar field (cxSM). The cxSM provides an interesting framework in view of the fact that no clear sign of new physics has been detected at the LHC up to now, as the cxSM can accommodate a SFOEWPT without predicting detectable traces at current and future runs of the LHC  Ashoorioon:2009nf. Consequently, the possibility of probing a SFOEWPT in the cxSM via GWs becomes particularly important. It is therefore crucial to make precise and accurate predictions for the parameters characterizing the transition and to ensure theoretical control over GW predictions. The model has been studied previously using the $4D$ approach in Refs. Chiang:2017nmu; Freitas:2021yng; Cho:2021itv; Zhang:2023mnu and the $3D$ approach in Ref. Schicho:2022wty. However, a comprehensive analysis of the parameter space facilitating a SFOEWPT and of the prospects for future GW detection comparing different perturbative methods has not yet been carried out. In this paper, we study thermodynamic observables of FOEWPTs and determine the parameter space regions that can potentially be probed with the LISA GW observatory [^1] The question whether a predicted GW signal is detectable or not at a particular experiment can ultimately only be answered if the theoretical uncertainties from the hydrodynamic modeling of GW production during the FOEWPT and experimental uncertainties, e.g. from unknown (astrophysical or cosmological) GW foregrounds are taken into account. These considerations lie beyond the scope of this work, where we focus on the theoretical uncertainties in the perturbative description of the EW phase transition., focusing especially on the theoretical uncertainties in the predictions and the impact of residual gauge dependence and renormalization- and matching-scale dependence. We compare and contrast the $4D$ and the $3D$ approaches, assessing their respective limitations in capturing the dynamics of the transition. Moreover, we test the validity of the $3D$ approach — which relies on the high-temperature expansion — by including higher-dimensional operators in the EFT Lagrangian.
@@ 3.4 High-temperature EFT
We use this subset of dimension-six operators to assess the validity of the EFT expansion. For this purpose, we do not take into account the gauge-dependence of these matching relations Chala:2025aiz, further operators at this order which include derivatives Chala:2024xll or corrections to the RGE running from the higher-dimensional operators Chala:2025cya. The impact of the operators in Eq.˜41 on thermodynamic observables is discussed in Section˜6. It should be noted that in the dimensionally reduced theory, scalar fields have mass dimension $1/2$, hence the Wilson coefficients in Eq.˜42 are dimensionless. However, we refer to them as dimension-six operators, as they correspond to operators with mass dimension six in 4D.
@@ 4.1.2 Gauge dependence
Next, we compare the different approaches defined at the beginning of this section ($\overline{\text{MS}}; 4D4D$-$\widetilde{\text{OS}}$; $3D$-$\overline{\text{MS}}$ at two-loop level with $\mathcal{O}(g^{4})$ matching) in terms of the numerical impact of the residual gauge-dependence on the predictions for the critical temperature $T_{c}$ and the latent heat $L_{c}$. This is depicted in Fig.˜9 for $T_{c}$ (left plot) and $L_{c}$ (right plot), with the solid lines showing the result for $\xi=0$, and the shaded bands indicate the theoretical uncertainty from the $\xi$-variation as discussed above. Moreover, the dashed lines show the predictions for the manifestly gauge-independent calculation, as determined using the $\hbar$-expansion about the tree-level minima of the effective potential (Eq.˜30). As already visible in Fig.˜8, we find good agreement between the gauge-independent results and the gauge-dependent results based on the direct potential minimization in the $3D$ case (blue). On the contrary, there is a significant deviation between the gauge-independent results based on the $\hbar$-expansion and the ones from direct numerical minimization of the loop-corrected potential in the $4D$ approach, irrespective of whether the $\overline{\text{MS}}$- or the $\widetilde{\text{OS}}-$prescription is used. The gauge-independent predictions for $T_{c}$ deviate by more than 50 is applied, and the gauge-independent predictions for $L_{c}/T_{c}^{4}\sim 2.5\text{--}4$ deviate by a factor of a few across the parameter space shown on the plot. The main culprit for these huge discrepancies in the $4D$ approaches lies in the fact that the tree-level minimum is significantly shifted by leading thermal loop corrections. Furthermore, in the $4D$ approach, the method to extract gauge-independent daisy resummation contributions, following Ref. Patel:2011th, is somewhat ad hoc, relying on evaluating different parts of the potential at different minima values. Taking into account the observations from the discussion of the renormalization scale dependence in Section˜3.2, our results suggest that the inherent theoretical uncertainties of the $4D$ approach at the one-loop level are neither captured by a scale variation about a factor of two, nor by the rather large $\xi$ variation that we have applied here. On the contrary, the proximity of the gauge-independent and the Landau gauge result in the $3D$ approach once again indicates a good perturbative behavior of the effective potential.
@@ 4.2 Nucleation thermodynamics
The percolation temperature $T_{p}$ is the temperature at which bubbles of the new phase have grown sufficiently to form a connected network that stretches over a significant fraction of the universe Guth:1981uk. At this point, the phase transition effectively completes, and we expect most of the GW produced during the transition to originate from this moment Ellis:2018mja, which occurs when the fraction of the extended volume occupied by the new phase is around 34 The extended volume double counts the overlapping regions of the bubbles, which is removed by exponentiating it Athron:2023xlk. The defining relation for $T_{p}$ is
@@ 7 Summary and conclusions
However, the $3D$ EFT approach is subject to an additional intrinsic theory uncertainty from truncating the EFT at finite operator dimension. To assess this uncertainty, we compared results including only renormalizable/dimension-four operators with those also incorporating dimension-six operators. At high temperatures, both types of results agree well, indicating a reliable behavior of the EFT. However, for lower temperatures, particularly below $T_{c}\lesssim 80$ GeV, the dimension-six operators noticeably affect the critical temperature and enhance the predicted strength of the phase transition and GW signals. We found that the impact on the critical temperature is more pronounced than on the percolation temperature. Nevertheless, the effects on the GW signals (which are computed at percolation) remain sizable. Specifically, we observed changes in the transition strength $\alpha$ of up to 100 up to an order of magnitude. Since stronger GW signals correlate with lower transition temperatures, these effects become relevant precisely where the prospects for a GW detection with LISA are best. Interestingly, the inclusion of dimension-six operators brings the predictions of the $3D$ approach closer to the results obtained using the $4D$-$\overline{\text{MS}}$ approach, which may be more reliable in this regime since it relies to a much lesser extent on the high-$T$ expansion. Still, since we included only a subset of all possible dimension-six operators, and since higher-dimensional terms are not considered with which the behavior of the EFT could be further checked, it remains unclear whether this signals a genuine improvement or merely a breakdown of the EFT. We leave a dedicated study in this direction for future work.
@@ Acknowledgements.
We thank Andreas Ekstedt, Guilherme Guedes, Panagiotis Stylianou and Johannes Braathen for useful discussions and comments. A.D., M.L. and G.W. acknowledge the support of the Deutsche Forschungsgemeinschaft (DFG, German Research Association) under Germany’s Excellence Strategy-EXC 2121 “Quantum Universe”-390833306. This work has also been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 491245950. T.B. acknowledges the support of the Spanish Agencia Estatal de Investigación through the grant “IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S”. The project that gave rise to these results received the support of a fellowship from the “la Caixa” Foundation (ID 100010434). The fellowship code is LCF/BQ/PI24/12040018. A.D. acknowledges the support by Consejería de Universidad, Investigación e Innovación, Gobierno de España and Unión Europea – NextGenerationEU under grant AST2_6.5.
@@ A.1 Input parameters
$\{m_{W},m_{Z},m_{h},m_{t}\}=\{80.379,\,91.1876,\,125,\,172.5\}\,\text{GeV},\quad G_{F}=1.1663787\times 10^{-5}\,\text{GeV}^{-2}.$ (73)
@@
$\frac{\delta g^{2}}{g^{2}}=\frac{\Sigma_{W}(m_{W}^{2})-\Sigma_{W}(0)}{m_{W}^{2}}-\frac{g^{2}}{16\pi^{2}}\left[4\ln\frac{\mu^{2}}{m_{W}^{2}}+\left(\frac{7}{2}\frac{m_{Z}^{2}}{m_{Z}^{2}-m_{W}^{2}}-2\right)\ln\frac{m_{W}^{2}}{m_{Z}^{2}}+6\right].$ (83)
@@
$\frac{\partial^{2}V_{1}^{T=0}}{\partial\phi_{s}^{2}}\,\bigg|_{\phi_{h}=v,\phi_{s}=0}$ $=\Sigma_{\phi_{s},\text{proper}}(0),$ (87a)
$\frac{\partial^{2}V_{1}^{T=0}}{\partial\phi_{h}^{2}}\,\bigg|_{\phi_{h}=v,\phi_{s}=0}$ $=\Sigma_{\phi_{h},\text{proper}}(0),$ (87b)
$\frac{\partial V_{1}^{T=0}}{\partial\phi_{h}}\,\bigg|_{\phi_{h}=v,\phi_{s}=0}$ $=T_{\phi_{h}},$ (87c)
@@
$\delta\lambda_{h}$ $=-\frac{2}{v^{2}}\left(\,\Sigma_{\phi,\text{proper}}(0)+\frac{1}{3}\Sigma_{\phi,\text{tad}}(0)\right)\,.$ (90c)

### v2

@@ 1 Introduction
In the Standard Model (SM), with the mass of the detected Higgs boson measured at about 125 GeV ATLAS:2012yve; CMS:2012qbp, electroweak (EW) symmetry breaking is predicted to proceed via a smooth crossover instead of a strong FOEWPT (SFOEWPT) Kajantie:1996mn. As a consequence, the presence of a SFOEWPT would be a clear indication of physics beyond the Standard Model (BSM). A SFOEWPT can occur even in simple extensions of the SM. For instance, models with additional scalar singlets Profumo:2007wc; Espinosa:2011ax; Cline:2012hg, SU(2) doublets Dorsch:2013wja; Basler:2016obg; Goncalves:2021egx; Biekotter:2021ysx; Biekotter:2022kgf; Biekotter:2023eil and/or triplets Patel:2012pi; Niemi:2018asa, effective theories incorporating higher dimensional operators Grojean:2004xa; Camargo-Molina:2021zgz; Camargo-Molina:2024sde, classically conformal models Nardini:2007me; Konstandin:2011dr; Kierkla:2022odc; Kierkla:2023von, or models with extra dimensions Creminelli:2001th can all exhibit a SFOEWPT. As mentioned above, a sufficiently strong FOEWPT can be probed by future space-based GW observatories. The observation of a GW signal would be very helpful in this context to pinpoint the underlying scalar sector responsible for the FOEWPT.
@@
Irrespective of the specific thermal resummation prescription, the main advantage of the $4D$ approach is that the form of the effective potential at one-loop level is relatively simple to implement, while capturing the leading thermal corrections and providing a qualitatively correct picture of the phase transitions in many extensions of the SM. Recent studies using the $4D$ approach have, for instance, been carried out for the singlet-extended SM Jiang:2015cwa; Chiang:2017nmu; Chiang:2018gsn; Carena:2019une; Cho:2021itv; Fernandez-Martinez:2022stj; Ellis:2022lft; Zhang:2023mnu; Braathen:2025svl, the Two Higgs doublet model Goncalves:2021egx; Biekotter:2022kgf; Biekotter:2021ysx and the triplet-extended SM Chala:2018opy; Zhou:2018zli, taking into account current experimental constraints. In some studies also two-loop contributions have been considered in the $4D$ approach Arnold:1992rz; Espinosa:1996qw; Laine:2000kv; Funakubo:2012qc; Bahl:2024ykv. A key drawback of the $4D$ approach is that it is often affected by the poor behavior of the perturbative expansion at finite temperature, leading to sizable renormalization scale and residual gauge dependencies in predictions for physical observables characterizing the phase transitions. In addition, it suffers from infrared divergences and an unsystematic treatment of thermal fluctuations in the evaluation of transition rates Ekstedt:2020abj; Athron:2022jyi. This leads to large theoretical uncertainties in the predictions for these parameters and quantities that are derived from them, such as GW signals produced during the transitions Croon:2020cgk; Lewicki:2024xan; Gould:2024jjt; Zhu:2025pht.
@@
Another problem occurring both in the $4D$ and the $3D$ approaches is the presence of residual gauge dependence in predictions for physical observables related to the phase transition. A proposed solution to this issue is the so-called $\hbar$-expansion Patel:2011th, which systematically expands physical quantities in powers of a loop-counting parameter $\hbar$. This ensures that one obtains gauge-independent results for physical observables order by order in perturbation theory. At the one-loop level in the $4D$ approach, applying the $\hbar$-expansion removes gauge dependence, but leads to results that, in many models, are not sufficiently precise, deviating substantially from more accurate predictions Patel:2011th. Additionally, the $\hbar$-expansion must be implemented consistently with the thermal resummation procedure to obtain a gauge-independent formulation at finite temperature Ekstedt:2018ftj; Athron:2022jyi; Qin:2024dfp.
@@
Specifically, we investigate the EW phase transition in the SM extended by a complex gauge singlet scalar field (cxSM). The cxSM provides an interesting framework in view of the fact that no clear sign of new physics has been detected at the LHC up to now, as the cxSM can accommodate a SFOEWPT without predicting detectable traces at current and future runs of the LHC  Ashoorioon:2009nf. Consequently, the possibility of probing a SFOEWPT in the cxSM via GWs becomes particularly important. It is therefore crucial to make precise and accurate predictions for the parameters characterizing the transition and to ensure theoretical control over GW predictions. The model has been studied previously using the $4D$ approach in Refs. Chiang:2017nmu; Freitas:2021yng; Cho:2021itv; Zhang:2023mnu; Ghosh:2025rbt; Jiang:2015cwa and the $3D$ approach in Ref. Schicho:2022wty. However, a comprehensive analysis of the parameter space facilitating a SFOEWPT and of the prospects for future GW detection comparing different perturbative methods has not yet been carried out. In this paper, we study thermodynamic observables of FOEWPTs and determine the parameter space regions that can potentially be probed with the LISA GW observatory [^1] The question whether a predicted GW signal is detectable or not at a particular experiment can ultimately only be answered if additional theoretical uncertainties, e.g. from the hydrodynamic modeling of GW production during the FOEWPT Caprini:2024hue and ambiguities in choosing an appropriate transition temperature for the evaluation Athron:2023rfq, as well as experimental uncertainties, e.g. from unknown (astrophysical or cosmological) GW foregrounds are taken into account. These considerations lie beyond the scope of this work, where we focus on the theoretical uncertainties in the perturbative description of the EW phase transition., focusing especially on the theoretical uncertainties in the predictions and the impact of residual gauge dependence and renormalization- and matching-scale dependence. We compare and contrast the $4D$ and the $3D$ approaches, assessing their respective limitations in capturing the dynamics of the transition. Moreover, we test the validity of the $3D$ approach — which relies on the high-temperature expansion — by including higher-dimensional operators in the EFT Lagrangian.
@@ 3.4 High-temperature EFT
We use this subset of dimension-six operators to assess the validity of the EFT expansion. For this purpose, we do not take into account the gauge-dependence of these matching relations Chala:2025aiz; Bernardo:2025vkz, further operators at this order which include derivatives Chala:2024xll; Chala:2025oul or corrections to the RGE running from the higher-dimensional operators Chala:2025cya. The impact of the operators in Eq.˜41 on thermodynamic observables is discussed in Section˜6. It should be noted that in the dimensionally reduced theory, scalar fields have mass dimension $1/2$, hence the Wilson coefficients in Eq.˜42 are dimensionless. However, we refer to them as dimension-six operators, as they correspond to operators with mass dimension six in 4D.
@@ 4.1.2 Gauge dependence
Next, we compare the different approaches defined at the beginning of this section ($\overline{\text{MS}}; 4D4D$-$\widetilde{\text{OS}}$; $3D$-$\overline{\text{MS}}$ at two-loop level with $\mathcal{O}(g^{4})$ matching) in terms of the numerical impact of the residual gauge-dependence on the predictions for the critical temperature $T_{c}$ and the latent heat $L_{c}$. This is depicted in Fig.˜9 for $T_{c}$ (left plot) and $L_{c}$ (right plot), with the solid lines showing the result for $\xi=0$, and the shaded bands indicate the theoretical uncertainty from the $\xi$-variation as discussed above. Moreover, the dashed lines show the predictions for the manifestly gauge-independent calculation, as determined using the $\hbar$-expansion about the tree-level minima of the effective potential (Eq.˜30). We note that in the $4D$-$\widetilde{\text{OS}}$ approach the finite counterterms are gauge-dependent (see Appendix˜A); hence, it is not possible to perform an $\hbar$-expansion leading to a gauge-independent prediction. As already visible in Fig.˜8, we find good agreement between the gauge-independent results and the gauge-dependent results based on the direct potential minimization in the $3D$ case (blue). On the contrary, there is a significant deviation between the gauge-independent results based on the $\hbar$-expansion and the ones from direct numerical minimization of the loop-corrected potential in the $4D$ approach, irrespective of whether the $\overline{\text{MS}}$- or the $\widetilde{\text{OS}}-$prescription is used. The gauge-independent predictions for $T_{c}$ deviate by more than 50% from the results where no $\hbar$-expansion is applied, and the gauge-independent predictions for $L_{c}/T_{c}^{4}\sim 2.5\text{--}4$ deviate by a factor of a few across the parameter space shown on the plot. The main culprit for these huge discrepancies in the $4D$ approaches lies in the fact that the tree-level minimum is significantly shifted by leading thermal loop corrections. Furthermore, in the $4D$ approach, the method to extract gauge-independent daisy resummation contributions, following Ref. Patel:2011th, is somewhat ad hoc, relying on evaluating different parts of the potential at different minima values. Taking into account the observations from the discussion of the renormalization scale dependence in Section˜3.2, our results suggest that the inherent theoretical uncertainties of the $4D$ approach at the one-loop level are neither captured by a scale variation about a factor of two, nor by the rather large $\xi$ variation that we have applied here. On the contrary, the proximity of the gauge-independent and the Landau gauge result in the $3D$ approach once again indicates a good perturbative behavior of the effective potential.
@@ 4.2 Nucleation thermodynamics
The percolation temperature $T_{p}$ is the temperature at which bubbles of the new phase have grown sufficiently to form a connected network that stretches over a significant fraction of the universe Guth:1981uk. At this point, the phase transition effectively completes, and we expect most of the GW produced during the transition to originate from this moment Ellis:2018mja; Athron:2023rfq, which occurs when the fraction of the extended volume occupied by the new phase is around 34%. [^22] This corresponds to the probability $P_{f}(T_{p})=e^{-0.34}\approx 71\%$ that a random point of space is still in the false vacuum. The extended volume double counts the overlapping regions of the bubbles, which is removed by exponentiating it Athron:2023xlk. The defining relation for $T_{p}$ is
@@ 7 Summary and conclusions
However, the $3D$ EFT approach is subject to an additional intrinsic theory uncertainty from truncating the EFT at finite operator dimension. To assess this uncertainty, we compared results including only renormalizable/dimension-four operators with those also incorporating dimension-six operators. At high temperatures, both types of results agree well, indicating a reliable behavior of the EFT. However, for lower temperatures, particularly below $T_{c}\lesssim 80$ GeV, the dimension-six operators noticeably affect the critical temperature and enhance the predicted strength of the phase transition and GW signals. We found that the impact on the critical temperature is more pronounced than on the percolation temperature. Nevertheless, the effects on the GW signals (which are computed at percolation) remain sizable. Specifically, we observed changes in the transition strength $\alpha$ of up to 100% and shifts in the GW peak amplitude by up to an order of magnitude. Since stronger GW signals correlate with lower transition temperatures, these effects become relevant precisely where the prospects for a GW detection with LISA are best. Interestingly, the inclusion of dimension-six operators brings the predictions of the $3D$ approach closer to the results obtained using the $4D$-$\overline{\text{MS}}$ approach, which may be more reliable in this regime since it relies to a much lesser extent on the high-$T$ expansion. Still, since we included only a subset of all possible dimension-six operators, and since higher-dimensional terms are not considered with which the behavior of the EFT could be further checked, it remains unclear whether this signals a genuine improvement or merely a breakdown of the EFT. We leave a dedicated study in this direction for future work.
@@ Acknowledgements.
We thank Andreas Ekstedt, Guilherme Guedes, Panagiotis Stylianou, Johannes Braathen, Philipp Schicho and Peter Athron for useful discussions and comments. A.D., M.L. and G.W. acknowledge the support of the Deutsche Forschungsgemeinschaft (DFG, German Research Association) under Germany’s Excellence Strategy-EXC 2121 “Quantum Universe”-390833306. This work has also been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 491245950. T.B. acknowledges the support of the Spanish Agencia Estatal de Investigación through the grant “IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S”. The project that gave rise to these results received the support of a fellowship from the “la Caixa” Foundation (ID 100010434). The fellowship code is LCF/BQ/PI24/12040018. A.D. acknowledges the support by Consejería de Universidad, Investigación e Innovación, Gobierno de España and Unión Europea – NextGenerationEU under grant AST2_6.5.
@@ A.1 Input parameters
$\{m_{W},m_{Z},m_{h},m_{t}\}=\{80.379,\,91.1876,\,125.1,\,172.5\}\,\text{GeV},\quad G_{F}=1.1663787\times 10^{-5}\,\text{GeV}^{-2}.$ (73)
@@
$\frac{\delta g^{2}}{g^{2}}=\frac{\Sigma_{W}^{(\overline{\text{MS}})}(m_{W}^{2})-\Sigma_{W}^{(\overline{\text{MS}})}(0)}{m_{W}^{2}}-\frac{g^{2}}{16\pi^{2}}\left[4\ln\frac{\mu^{2}}{m_{W}^{2}}+\left(\frac{7}{2}\frac{m_{Z}^{2}}{m_{Z}^{2}-m_{W}^{2}}-2\right)\ln\frac{m_{W}^{2}}{m_{Z}^{2}}+6\right].$ (83)
@@
$\frac{\partial^{2}V_{1}^{T=0}}{\partial\phi_{s}^{2}}\,\bigg|_{\phi_{h}=v,\phi_{s}=0}$ $=\Sigma_{\phi_{s},\text{proper}}^{(\overline{\text{MS}})}(0),$ (87a)
$\frac{\partial^{2}V_{1}^{T=0}}{\partial\phi_{h}^{2}}\,\bigg|_{\phi_{h}=v,\phi_{s}=0}$ $=\Sigma_{\phi_{h},\text{proper}}^{(\overline{\text{MS}})}(0),$ (87b)
$\frac{\partial V_{1}^{T=0}}{\partial\phi_{h}}\,\bigg|_{\phi_{h}=v,\phi_{s}=0}$ $=T_{\phi_{h}}^{(\overline{\text{MS}})},$ (87c)
@@
$\delta\lambda_{h}$ $=-\frac{2}{v^{2}}\left(\,\Sigma_{\phi,\text{proper}}^{(\overline{\text{MS}})}(0)+\frac{1}{3}\Sigma_{\phi,\text{tad}}^{(\overline{\text{MS}})}(0)\right)\,.$ (90c)